The Fraunhofer diffraction pattern due to a single slit of width 0.4 cm is obtained

with the help of a lens of focal length 30 cm. If the wavelength of light used is

5890 Å, calculate the distance of the first dark fringe and the consecutive bright

fringe from the axis.


1
Expert's answer
2021-05-04T06:43:40-0400

Given, "D=30cm, d=0.4cm ,\lambda=5890A=5.89\times 10^{-7}m"


As we know, "x_n=\dfrac{nD\lambda}{d}=\dfrac{1\times 30\times 5.89\times 10^{-7}}{0.4}=\dfrac{176.7\times 10^{-7}}{0.4}"


"=441\times .78\times 10^{-7}=0.0441mm"